Kleen Company acquired patent rights on January 10 of Year 1 for $882,000. The patent has a useful life equal to its legal life of eight years. On January 7 of Year 4, Kleen successfully defended the patent in a lawsuit at a cost of $39,600. Required: a. Determine the patent amortization expense for Year 4 ended December 31. b. Journalize the adjusting entry on December 31 of Year 4 to recognize the amortization. Refer to the Chart of Accounts for exact wording of account titles.
Answer #1

a) Patent amortization expense for year 4 :

Original amortization = 882000/8 = 110250 per year

Amortization for year 4 = (882000-330750+39600)/5 = 118170

b) Journal entries

Date account and explanation debit credit
Dec 31, Year 4 Amortization expense 118170
Patent 118170
(To record amortization)